# m2mKD: Module-to-Module Knowledge Distillation for Modular Transformers

### v1

@@ 6 Conclusion and Future works
In this work, we present module-to-module knowledge distillation (m2mKD) as a general approach for transferring knowledge in the context of modular model training. m2mKD leverages a monolithic teacher model to facilitate the training of a modular student model, offering a promising strategy for transforming pretrained monolithic models into modular architectures to harness their advantages. Experimental results on NACs and MoEs demonstrate that the proposed pipeline effectively enhances the IID accuracy and OOD robustness of modular models, even in scenarios where the student model size is smaller than the teacher model size. However, there is still room for improvement in bridging the performance gap between the student modular model and the teacher model. To further enhance the effectiveness, m2mKD can be combined with existing techniques such as ensemble learning. We hope that this work will advance research on module-wise knowledge distillation and monolithic-to-modular conversion, and we look forward to future studies in these areas.
Appendix A Notations
Table 8: Notations.
@@ Appendix A Notations
Appendix B Training Hyperparameters
Table 9: Hyperparameters for the E2E training phase of NACs.
@@ Appendix B Training Hyperparameters
Table 10: Hyperparameters for the few-shot training of NACs, DeiT and V-MoE-B models.
NACs DeiT / V-MoE-B
@@
Appendix C Ablation studies
@@
In the proposed m2mKD pipeline, the stitch layers are used only during the distillation phase and subsequently discarded to ensure no modifications are required for the student model. Although the stitch layers typically account for just a small proportion of the total parameters in a large modular model (e.g., 1% for our V-MoE-B student), directly removing them may result in the loss of some knowledge acquired from the teacher, potentially impacting the final model’s performance negatively. To investigate this, we conduct experiments preserving the stitch layers in the student model. However, as shown in Table 11, the results indicate no noticeable improvement from this approach. Consequently, we maintain our decision to discard the stitch layers.
@@
Some additional experiments are conducted based on the setting of preserving stitch layers. The previous experiment indicates that the presence of the stitch layers does not influence much, so we suppose the following conclusions are applicable regardless of their removal. Firstly, since the NAC models are able to achieve great improvement with student modules distilled for merely 10 epochs, we test the performance of V-MoE-B models under the same condition. To this end, we adjust some of the hyperparameters of the m2mKD phase and distill the student module for 10 epochs. The E2E phase remains to be 100 epochs. Table 11 shows that decreasing the distillation epochs from 100 to 10, the accuracy drops 2.1%, but is still 2.8% better than the Deep Incubation baseline. We apply the same changes on the incubation phase of Deep Incubation and, surprisingly, there is no degradation of the performance using the new hyperparameters. Next, we inspect the scalability of m2mKD. The DeiT-Huge model (632M) pretrained by Deep Incubation is chosen as the teacher. The student model ($L=4$) is a V-MoE-Large model with 12 MoE layers placed on every other layer, which totally contains 1.0B parameters. Table 11 compares m2mKD with Deep Incubation and the former achieves 0.27% higher accuracy on ImageNet-1k.
@@
The remaining ablation experiments are conducted based on the preservation of stitch layers. The previous experiment indicates that the presence of the stitch layers does not have a significant influence, we believe that the following conclusions remain valid regardless of their removal. Firstly, since the NAC models demonstrate considerable improvement with student modules distilled for only 10 epochs, we test the performance of V-MoE-B models under the same condition. To this end, we adjust some of the hyperparameters during the m2mKD phase and distill the student module for 10 epochs, while retaining the E2E phase at 100 epochs. As shown in Table 11, decreasing the distillation epochs from 100 to 10 results in a drop of 2.1% in accuracy, yet it remains 2.8% higher than the Deep Incubation baseline. Surprisingly, applying the same changes to the incubation phase of Deep Incubation does not lead to any performance degradation. Next, we examine the scalability of m2mKD. The teacher model is the DeiT-Huge model (632M) pretrained by Deep Incubation, while the student model ($L=4$) is a V-MoE-Large model with 12 MoE layers placed on every other layer, resulting in a total of 1.0B parameters. As presented in Table 11, m2mKD outperforms Deep Incubation by 0.27% in terms of accuracy on ImageNet-1k.

### v2

@@ 6 Conclusion and Future Works
In this work, we present module-to-module knowledge distillation (m2mKD) as a general approach for transferring knowledge in modular model training. m2mKD leverages a monolithic teacher model to facilitate the training of a modular student model, offering a promising strategy for transforming pretrained monolithic models into modular architectures to harness their advantages. Experimental results on NACs and MoEs demonstrate that the proposed pipeline enhances the IID accuracy and OOD robustness, even when the student model size exceeds that of the teacher model. However, there is still room for narrowing the performance gap between the student and teacher. To further enhance effectiveness, combining m2mKD with techniques like ensemble learning holds potential. We hope that this work can advance research on module-wise knowledge distillation and monolithic-to-modular conversion.
Appendix 0.A Notations
Table 4: Notations commonly used in the main paper.
@@ Appendix 0.A Notations
Appendix 0.B Training Hyperparameters
Table 5: A selection of hperparameters for the E2E training phase of NACs. When training on the same dataset, the hyperparameters remain unchanged for both the reproduced baselines and the distilled NAC models.
@@ Appendix 0.B Training Hyperparameters
Table 6: Hyperparameters for the few-shot training of NAC, DeiT and V-MoE-B models.
NAC DeiT / V-MoE-B
@@
Appendix 0.C Reproducibility
IID OOD
(Tiny-ImageNet) (Tiny-ImageNet-R)
Acc@1 Acc@5 Acc@1 Acc@5
1st 66.47 85.08 24.31 44.38
2nd 66.13 85.33 24.84 45.00
3rd 65.94 85.49 25.29 45.61
MEAN 66.18 85.30 24.81 45.00
STDEV 0.27 0.21 0.49 0.62
Appendix 0.D Few-shot results
Seed 1-shot 2-shot 4-shot 8-shot
E2E m2mKD E2E m2mKD E2E m2mKD E2E m2mKD
1 53.66 60.82 64.39 64.39 71.21 75.06 81.03 80.53
2 60.82 50.08 64.39 59.03 75.06 78.91 84.04 82.03
3 67.97 67.97 75.13 67.97 78.91 79.87 80.53 79.03
4 60.82 60.82 67.97 57.24 75.06 78.91 84.04 82.53
5 50.08 60.82 60.82 60.82 73.14 78.91 80.03 79.53
MEAN 58.67 60.10 66.54 61.89 74.68 78.33 81.93 80.73
STDEV 6.98 6.40 5.43 4.31 2.85 1.88 1.95 1.52
Table 9: Differences between two runs of the 2-shot experiment for NAC models.
seed E2E m2mKD
1st 2nd $|\Delta|$ 1st 2nd $|\Delta|$
1 64.39 69.76 5.37 64.39 66.18 1.79
2 64.39 64.39 0.00 59.03 60.82 1.79
3 75.13 75.13 0.00 67.97 75.13 7.16
4 67.97 66.18 1.79 57.24 62.60 5.36
5 60.82 62.60 1.78 60.82 64.39 3.57
MEAN 66.54 67.61 1.07 61.89 65.82 3.93
STDEV 5.43 4.97 0.46 4.31 5.57 1.26
Table 10: Few-shot performance of DeiT and V-MoE-B models on CIFAR-100 (top) and CUB-2011 (bottom). The models are tasked with classifying samples into 8 classes (i.e., 8-way). We compare three models: a DeiT-B model trained using Deep Incubation (DeiT-DI), and two V-MoE-B models trained using Deep Incubation (MoE-DI) and m2mKD (MoE-Ours), respectively.
CIFAR-100
Seed 1-shot 5-shot 10-shot
DeiT-DI MoE-DI MoE-Ours DeiT-DI MoE-DI MoE-Ours DeiT-DI MoE-DI MoE-Ours
1 70.83 70.83 70.83 90.83 90.00 89.17 90.83 91.67 90.42
2 70.83 62.50 66.37 90.00 87.50 86.67 92.92 92.50 90.42
3 87.50 95.83 91.67 95.83 94.17 94.17 97.08 96.25 96.67
4 95.83 79.17 87.50 95.83 94.17 94.17 97.50 94.58 96.25
5 79.17 70.83 70.83 88.33 87.50 87.50 90.00 89.58 90.42
MEAN 80.83 75.83 77.44 92.16 90.67 90.34 93.67 92.92 92.84
STDEV 10.87 12.64 11.33 3.47 3.36 3.61 3.48 2.59 3.31
CUB-2011
Seed 1-shot 5-shot 10-shot
DeiT-DI MoE-DI MoE-Ours DeiT-DI MoE-DI MoE-Ours DeiT-DI MoE-DI MoE-Ours
1 83.33 70.83 70.83 90.83 90.83 91.67 95.00 94.38 93.75
2 79.17 83.33 87.5 94.17 93.33 94.17 98.75 95.63 98.13
3 95.83 91.67 95.83 95.00 92.50 92.50 94.38 91.25 91.25
4 87.50 75.00 79.17 88.33 80.83 84.17 84.38 80.63 86.25
5 66.67 70.83 75.00 90.83 88.33 94.17 96.25 93.75 95.63
MEAN 82.50 78.33 81.67 91.83 89.16 91.34 93.75 91.13 93.00
STDEV 10.78 9.04 10.03 2.73 5.04 4.15 5.50 6.08 4.54
Appendix 0.E Ablation Studies
@@
In the proposed m2mKD pipeline, stitch layers are used only during distillation and subsequently discarded to avoid modifying the student model. Although they typically account for a small fraction of parameters in large modular models (e.g., 1% for our V-MoE-B student), their removal may lead to loss of knowledge acquired from the teacher, potentially impacting performance. To investigate this aspect, we conduct experiments preserving stitch layers in the student model. However, as shown in Tab. 11, the results indicate no noticeable improvement from this approach. Consequently, we maintain our decision to discard the stitch layers.
@@
The remaining ablation experiments are conducted based on the preservation of stitch layers. Given the negligible influence of preserving stitch layers, we believe that the following conclusions remain valid even if they are removed. Firstly, since the NAC models demonstrate considerable improvement with student modules distilled for only 10 epochs, we test the performance of V-MoE-B models under the same condition. To this end, we adjust some of the hyperparameters during the m2mKD phase and distill the student module for 10 epochs, while retaining the E2E phase at 100 epochs. As shown in Tab. 11, decreasing the distillation epochs from 100 to 10 results in a drop of 2.1% in accuracy, yet it remains 2.8% higher than the Deep Incubation baseline. Surprisingly, applying the same changes to the incubation phase of Deep Incubation does not lead to any performance degradation. Next, we examine the scalability of m2mKD. The teacher model is the DeiT-Huge model (632M) pretrained by Deep Incubation, while the student model ($L=4$) is a V-MoE-Large model with 12 MoE layers placed on every other layer, resulting in a total of 1.0B parameters. As presented in Tab. 11, m2mKD outperforms Deep Incubation by 0.27% in terms of accuracy on ImageNet-1k.